OpenStudy (anonymous):

Solve the equation: (4x − 6)(2x − 5) = 0

Parth (parthkohli):

Do you want me to teach this to you?

mathslover (mathslover):

k parth u carry on

mathslover (mathslover):

@jiteshmeghwal9 please dont answer

OpenStudy (jiteshmeghwal9):

i mn't giving answer

mathslover (mathslover):

ok .. parth u carry on with ur teaching

OpenStudy (jiteshmeghwal9):

i wanna let him understnd

Parth (parthkohli):

Okay, @herbdog You need to be familiar with 'zero-product-rule' okay? The name is intimidating, but it's understandable.

Parth (parthkohli):

Have you heard that if you multiply 0 to a number, then you get 0?

OpenStudy (anonymous):

yes

OpenStudy (jiteshmeghwal9):

now let me teach also

Parth (parthkohli):

So, maybe 2x - 5 is 0.. that may be the reason why we have 0.

Parth (parthkohli):

Or maybe 4x - 6 is 0.

Parth (parthkohli):

Get it till now? @herb

OpenStudy (anonymous):

yes

Parth (parthkohli):

So we can have two \(x\)'s here, because we have two numbers that makes this zero.

Parth (parthkohli):

Can we now say that either \(4x - 6 = 0\) or \(2x - 5 = 0\)?

OpenStudy (anonymous):

yeah

Parth (parthkohli):

So as I taught you - can you solve these two equations?\[4x - 6 = 0\\2x - 5=0 \]These are two different equations, all right?

OpenStudy (anonymous):

okay, but cant we subtract the two equations or no?

OpenStudy (jiteshmeghwal9):

If given a polynomial, we sometimes want to know for which value of 'x' it is zero . For example: we solved linear equations of the type 2x+3=0 . we find that x=-3/2 is a solution to this equation. another way to say the same thing is that -3/2 is a zero of the polynomial. & remember it depends on the equation that how many zeros are founded.

Parth (parthkohli):

No, herb. They are not the same equations. Two different equations that you have to solve differently.

OpenStudy (anonymous):

oh, so if they different how do we solve it

Parth (parthkohli):

First solve this equation for me:\[4x - 6 = 0 \]

OpenStudy (anonymous):

isnt it -2x=0

Parth (parthkohli):

Nope. Add 6 to both sides.

OpenStudy (anonymous):

ooh becomes 4x=6

Parth (parthkohli):

Yeah, now divide both sides by 4.

OpenStudy (anonymous):

becomes 1.5

Parth (parthkohli):

Right - there's one solution. You have to get another solution by solving\[2x - 5 = 0 \]

OpenStudy (anonymous):

so i add 5 to both sides becomes 2x=5 right?

Parth (parthkohli):

Yes! You getting it! Now divide both by 2.

OpenStudy (anonymous):

and i get 2.5

Parth (parthkohli):

Woo! There you go! The two solutions are 1.5 and 2.5!

OpenStudy (anonymous):

so x = 1.5 & 2.5

Parth (parthkohli):

That's it :)

OpenStudy (anonymous):

thanks that helped alot

Parth (parthkohli):

I bet it did. You're welcome :)